OpenStudy (anonymous):

I need immediate help!!!!!!!! evaluate lim (as x goes to infinity) 2^x-e^10^x/3^x+5^x

OpenStudy (anonymous):

no 

OpenStudy (anonymous):

is that 10^x, or the whole thing?

OpenStudy (anonymous):

(a^m)^n = a^(mn), it's a basic exponent rule : (2^2)^3 = 2^6

OpenStudy (anonymous):

whole thing

OpenStudy (anonymous):

is it 2^x-(e^10^x/3^x)+5^x or (2^x-e^10^x)/(3^x+5^x)?

OpenStudy (anonymous):

the latter is true

OpenStudy (anonymous):